# Scaling of diffusion constants in perturbed easy-axis Heisenberg spin chains

### v1

@@ Abstract
Introduction. The study of nonequilibrium processes in quantum many-body systems continues to be one of the central endeavors in different fields of modern physics [1, 2, 3, 4, 5], ranging from fundamental questions in statistical mechanics to applied questions in material science. In this context, transport is a paradigmatic example of a nonequilibrium process with relevance to closed and open quantum systems alike. The understanding of nonequilibrium physics in general and transport in particular has seen substantial progress [6], due to experimental advances, fresh theoretical concepts, and the development of sophisticated analytical techniques and numerical methods.
@@
Since integrable systems are rather the exception than the rule, nonintegrable systems are the generic situation in physics. While these systems are expected to typically exhibit diffusive transport, the precise value of diffusion constants constitutes a highly nontrivial question and is a challenge for theory [21, 22, 23, 24, 25, 26, 27, 28, 29]. An equally nontrivial question is the dynamical behavior in the close vicinity of integrability [30, 31, 32, 33, 34, 35, 36]. This question includes the stability of superdiffusion at the isotropic point [37, 38, 39, 40, 41], but also the change of diffusion constants in the easy-axis regime [42, 43], where this change might be discontinuous. While such a change is in line with breaking microscopic integrability, it is still unexpected from macroscopic phenomenology, where a small perturbation of a normal conductor does not cause a significant change of the transport coefficient.
@@
Model and methods. In our work, a central model is the spin-$1/2$ XXZ chain [6],
@@
where the $S_{r}^{j}$ $(j=x,y,z)$ are spin-$1/2$ operators at site $r$, $N$ is the number of sites, $J>0$ is the antiferromagnetic exchange coupling constant, and $\Delta$ is the anisotropy in the $z$ direction. For all values of $\Delta$, the model in Eq. (1) is integrable and the total magnetization $S^{z}=\sum_{r}S_{r}^{z}$ is conserved, $[S^{z},H]=0$. Throughout our work, we focus on the easy-axis regime $\Delta>1$ and choose the specific value $\Delta=1.5$. It is important to note that we employ periodic boundary conditions, $S_{N+1}^{j}=S_{1}^{j}$, in the closed system but also in the open system introduced later.
@@
In our work, we apply dynamical quantum typicality for the numerical calculation of the quantity in Eq. (6), which allows us to treat comparatively large system sizes outside the range of standard exact diagonalization [46, 47]. This approach to finite systems is complemented by another technique, which is designed to numerically calculate $D$ directly in the thermodynamical limit. This technique is a recently proposed recursion method (RM) [27] and is based on Lanczos coefficients in Liouville space. Details of this method can be found in the supplemental material [48].
Figure 1: Diffusion constant $D$ versus perturbation strength $\Delta^{\prime}$ for the model $H_{\text{NNN}}$ in Eq. (2) with anisotropy $\Delta=1.5$, as calculated numerically by dynamical quantum typicality for different system sizes $N$ and by the recursion method in the thermodynamic limit [49]. (a) $\langle S^{z}\rangle=0$, $tJ=10$. (b) $\langle S^{z}\rangle=0$, $tJ=40$. (c) $S^{z}=0$, $tJ=10$. (d) $S^{z}=0$, $tJ=40$. In (d), existing results from the microcanonical Lanczos method are also depicted [43].
@@
[…]
@@
Figure 3: (a) Diffusion constant $D$ versus perturbation strength $\Delta^{\prime}$ for the model $H_{\text{NNN}}$ in Eq. (2) with anisotropy $\Delta=1.5$, as calculated numerically in the open system for different system sizes $N$. A comparison to the closed-system results from the recursion method is also indicated. (b) Extrapolation to the thermodynamic limit based on the linear scaling in (c).
A practical advantage of the Lindblad equation is that it can be solved by sophisticated numerical techniques, like stochastic unraveling [51, 15] or simulations based on matrix product states (MPS) [52, 53, 54]. In particular, MPS simulations are feasible for quite large system sizes. Yet, they are restricted to open boundary conditions and to large values of the system-bath coupling $\gamma$, due to the required reduction of entanglement growth. Thus, we use as an alternative strategy a recently proposed approach [55, 56, 57]. On the one hand, this approach is suitable for periodic boundary conditions and small system-bath coupling $\gamma$ [58]. On the other hand, it allows us to predict the open-system dynamics just on the knowledge of spatio-temporal correlation functions in the closed systems, which in turn enables us to treat open systems outside the range of exact diagonalization and stochastic unraveling. Details on this approach can be found in the supplemental material [48].
Results. We now move on to our results, starting with the closed-system scenario and the model in Eq. (2). We first focus on results from the recursion method, which are, as discussed in the supplemental material, well converged w.r.t. to the number of Lanczos coefficients and depicted in Fig. 1 (a). As clearly visible, the diffusion constant $D$ has a continuous dependence on the perturbation strength $\Delta^{\prime}$, with no obvious signature of a discontinuity in the weak-perturbation limit $\Delta^{\prime}\to 0$. Furthermore, for the values of $\Delta^{\prime}$ considered, $D(\Delta^{\prime})$ has a minor dependence on $\Delta^{\prime}$ and is an almost flat curve.
In addition, we compare in Fig. 1 (a) the results from the recursion method to finite-size results for the quantity in Eq. (6) from dynamical quantum typicality, for which raw data can be found in the supplemental material. We do so for the grandcanonical ensemble $\langle S^{z}\rangle=0$ and at a time $tJ=10$. Apparently, the finite-size data is in very good agreement with the recursion method and shows no significant dependence on system size. When redoing the comparison for a longer time $tJ=40$ in Fig. 1 (b), the agreement becomes worse, due to finite-size effects at nonweak perturbation strengths $\Delta^{\prime}$. Still, when the system size is increased, the finite-size data gets closer and closer to the result from the recursion method, and they might eventually coincide in the thermodynamic limit.
@@
Simple model. To understand this observation, let us for a moment come back to the closed system and use a simple model for the quantity $D(t)$ in Eq. (6), as shown in Fig. 4. After an initial linear increase, $D(t)$ has a plateau but, after a finite-size time $\propto N$, increases linearly with a finite-size slope $\propto 1/N$. This overall behavior describes well the unperturbed case $\Delta^{\prime}=0$ with periodic boundary conditions [59]. It is worth mentioning that the linear slope is the finite-size Drude weight [6]. Using this model for $D(t)$, we can calculate the spreading of an excitation,
@@
Assuming that $t^{*}$ is the relevant time scale for the open system, we plug in $t^{*}$ in the original model for $D(t)$ and use $D(t^{*})$ as an prediction for the diffusion constant in the open system. In Fig. 4 (b), we depict the finite-size scaling of the so obtained $D(t^{*})$, which is, despite the simplicity of arguments, very similar to the one for the open-system diffusion constant in Fig. 3 (c). Importantly, however, $D(t^{*})$ departs from a $1/N$-scaling for large $N$ and eventually diverges in the limit $N\to\infty$. The departure results from the fact that the steady-state time $t^{*}$ increases faster with $N$ than the time at which $D(t)$ is converged w.r.t. system size. This finding opens the possibility of a diverging diffusion constant in the unperturbed open system with periodic boundary conditions. Additional results for open boundary conditions can be found in the supplemental material [48].
Conclusion. We have analyzed the scaling of diffusion constants under integrability-breaking perturbations of the spin-$1/2$ XXZ chain in the easy-axis regime. To this end, we have used a variety of methods and covered both, closed and open systems, focusing on periodic boundary conditions. In the closed system, we have found evidence for a continuous change of diffusion constants over the full range of perturbation strengths. In the open system, we have found diffusion constants in quantitative agreement with the ones in the closed system in a range of nonweak perturbations. In the weak-perturbation limit, however, we have found disagreement and also pointed out the possibility of a diverging diffusion constant in the open system.
Acknowledgments. We thank Jochen Gemmer, Jacek Herbrych, Zala Lenarčič, Marcin Mierzejewski, and Peter Prelovšek for fruitful discussions.
@@
Supplemental material of
“Scaling of diffusion constants in perturbed easy-axis Heisenberg spin chains”
Calculation of equilibrium transport
Figure S1: Diffusion constant $D_{R}$ versus $R$, as obtained from the recursion method for the model (a) $H_{\text{NNN}}$ in Eq. (2) and (b) $H_{\text{B}}$ in Eq. (3).
Recursion method
@@
${F}(s)=\int_{0}^{\infty}\text{d}t\,e^{ts}\,\tilde{C}(t)$ (S1)
@@
$\tilde{C}(t)=\langle\tilde{j}(t)\tilde{j}\rangle\,,\quad\tilde{j}=\frac{j}{\sqrt{\langle j|j\rangle}}$ (S2)
of the normalized current. In particular, ${\cal F}(0)$ determines the diffusion constant $D=\langle j^{2}\rangle F(0)/\chi$. According to the Mori formalism, Eq. (S1) can be written in the form
${F}(s)=\frac{1}{s+\frac{b_{1}^{2}}{s+\frac{b_{2}^{2}}{s+\frac{b_{3}^{2}}{\cdots}}}}\,,$ (S3)
@@
$D=\frac{\langle j^{2}\rangle}{\chi b_{1}}\prod_{n=1}^{\infty}\left(\frac{b_{n}}{b_{n+1}}\right)^{(-1)^{n}}\,,$ (S4)
@@
$(O_{m}|O_{n})=\text{tr}[O^{\dagger}_{m}O_{n}]\,,$ (S5)
which defines a norm via
$||O||=\frac{\sqrt{(O|O)}}{\text{tr}[1]}\,.$ (S6)
@@
$\langle O(t)O\rangle\propto C(t)=\frac{(O|e^{i{\cal L}t}|O)}{||O||^{2}}\,.$ (S7)
Figure S2: Decay of the current autocorrelation function $\langle j(t)j\rangle$ in the unperturbed model in Eq. (1) with anisotropy $\Delta=1.5$ for (a) the grandcanonical ensemble $\langle S^{z}\rangle=0$ and (c) the canonical ensemble $S^{z}=0$, which is numerically calculated for different system sizes $N$ by the use of dynamical quantum typicality. (b) and (d) Corresponding time dependence of the diffusion coefficients $D(t)$.
Figure S3: The same data as in Fig. S2 but now for the model in Eq. (2) with $\Delta^{\prime}=0.4$.
@@
$b_{n}=||O^{\prime}_{n}||\,,\quad|O_{n})=\frac{|O^{\prime}_{n})}{b_{n}}\,.$ (S10)
@@
$b_{n}=\alpha n+\beta\,,\quad n\geq R\,,$ (S11)
@@
$F_{0}\approx D_{R}=\begin{cases}\frac{1}{\tilde{p}_{R}b_{R}}\prod_{m=1}^{\frac{R}{2}}\frac{b_{2m}^{2}}{b_{2m-1}^{2}}&\text{, even }R\\ \frac{\tilde{p}_{R}}{b_{R}}\prod_{m=1}^{\frac{R-1}{2}}\frac{b_{2m}^{2}}{b_{2m-1}^{2}}&\text{, odd }R\end{cases}\,,$ (S12)
@@
$p_{R}=\frac{\Gamma(\frac{R}{2}+\frac{\beta}{2\alpha})\Gamma(\frac{R}{2}+\frac{\beta}{2\alpha}+1)}{\Gamma^{2}(\frac{R}{2}+\frac{\beta}{2\alpha}+\frac{1}{2})}\,.$ (S13)
@@
$\alpha_{R}=b_{R}-b_{R-1}\,,\quad\beta_{R}=Rb_{R-1}-(R-1)b_{R}\,.$ (S14)
In Fig. S1, we summarize the so obtained $D_{R}$. As mentioned in the main text, $D_{R}$ is well converged w.r.t. $R$, for the two models and the perturbation strengths considered.
.1 Dynamical quantum typicality
@@
$\langle j(t)j\rangle=\frac{\langle\psi(t)|j|\varphi(t)\rangle}{\langle\psi(0)|\psi(0)\rangle}+{\cal O}\left(\frac{1}{\sqrt{d}}\right)\,,$ (S15)
@@
$\psi(t)=e^{-iHt}\,|\phi\rangle\,,\quad\varphi(t)=e^{-iHt}j\,|\phi\rangle\,,$ (S16)
and $|\phi\rangle$ is a Haar-random pure state. Importantly, the time argument in the approximation in Eq. (S15) is now a property of the two pure states and not of the current operator anymore. These pure states can be propagated in time using fourth-order Runge-Kutta or Chebyshev polynomials, which then gives access to comparatively large system sizes outside the range of standard exact diagonalization. An approximation similar to the one in Eq. (S15) can be also obtained for other operators, which we employ for the spatial-temporal correlation functions discussed later in the context of open systems.
In Figs. S2 - S4, we depict raw data for the current autocorrelation function $\langle j(t)j\rangle$ and its time integral $D(t)$ for three examples: The integrable model in Eq. (1) and the nonintegrable models in Eqs. (2) [with $\Delta^{\prime}=0.4$] and (3) [with $B/J=0.3$]. We do so for the two ensembles, grandcanonical $\langle S^{z}\rangle=0$ and canonical $S^{z}=0$. This and similar raw data is the basis of our analysis of diffusion constants in the main text. It is worth mentioning that the finite-size effects are unrelated to the approximation error in Eq. (S15), which is negligibly small in all cases considered.
In Figs. S5 and S6, we additionally depict for the two nonintegrable cases the current autocorrelation function in frequency space, as obtained from a Fourier transform up to a maximum time $T_{\text{max}}$. As can be seen, the specific choice of $T_{\text{max}}$ controls the frequency resolution, and a reasonable choice in a finite system is not obvious. As an additional consistency check of our approach, we compare to existing data from the microcanonical Lanzos method [43] and find a convincing agreement, as visible in Figs. S5 (c) and S6 (d). MCLM data at zero frequency is also shown in the main text.
Figure S4: The same data as in Fig. S2 but now for the model in Eq. (3) with $B/J=0.3$.
Calculation of nonequilibrium transport
Time evolution of densities
Figure S5: Diffusion coefficient $D$ versus frequency $\omega$ for the model in Eq. (2) with $\Delta^{\prime}=0.4$, as obtained numerically for the grandcanonical ensemble $\langle S^{z}\rangle=0$ using (a) fixed system size $N=24$ and different times $t_{\text{max}}$ and (b) fixed $t_{\text{max}}$ and different $N$. (c) and (d) show (a) and (b) for the canonical ensemble $S^{z}=0$. In (c), the MCLM result from Ref. [43] is also depicted.
Figure S6: The same data as in Fig. S5 but now for the model in Eq. (3) with $B/J=0.3$.

### v2

@@ Abstract
I Introduction
The study of nonequilibrium processes in quantum many-body systems continues to be one of the central endeavors in different fields of modern physics [1, 2, 3, 4, 5], ranging from fundamental questions in statistical mechanics to applied questions in material science. In this context, transport is a paradigmatic example of a nonequilibrium process with relevance to closed and open quantum systems alike. The understanding of nonequilibrium physics in general and transport in particular has seen substantial progress [6], due to experimental advances, fresh theoretical concepts, and the development of sophisticated analytical techniques and numerical methods.
@@ I Introduction
Since integrable systems are rather the exception than the rule, nonintegrable systems are the generic situation in physics. While these systems are expected to typically exhibit diffusive transport, the precise value of diffusion constants constitutes a highly nontrivial question and is a challenge for theory [21, 22, 23, 24, 25, 26, 27, 28]. An equally nontrivial question is the dynamical behavior in the close vicinity of integrability [29, 30, 31, 32, 33, 34, 35]. This question includes the stability of superdiffusion at the isotropic point [36, 37, 38, 39, 40], but also the change of diffusion constants in the easy-axis regime [41, 42, 43], where this change might be discontinuous. While such a change is in line with breaking microscopic integrability, it is still unexpected from macroscopic phenomenology, where a small perturbation of a normal conductor does not cause a significant change of the transport coefficient.
@@
II Model and methods
In our work, a central model is the spin-$1/2$ XXZ chain [6],
@@ II Model and methods
where the $S_{r}^{j}$ $(j=x,y,z)$ are spin-$1/2$ operators at site $r$, NN is the number of sites, $J>0$ is the antiferromagnetic exchange coupling constant, and $\Delta$ is the anisotropy in the $z$ direction. For all values of $\Delta$, the model in Eq. (1) is integrable and the total magnetization $S^{z}=\sum_{r}S_{r}^{z}$ is conserved, $S^{z},H]=0$. Throughout our work, we focus on the easy-axis regime $\Delta>1$ and choose the specific value $\Delta=1.5$. It is important to note that we employ periodic boundary conditions, $S_{N+1}^{j}=S_{1}^{j}$, in the closed system but also in the open system introduced later.
@@
In our work, we apply dynamical quantum typicality for the numerical calculation of the quantity in Eq. (6), which allows us to treat comparatively large system sizes outside the range of standard exact diagonalization [46, 47]. This approach to finite systems is complemented by another technique, which is designed to numerically calculate $D$ directly in the thermodynamical limit. This technique is a recently proposed recursion method (RM) [26] and is based on Lanczos coefficients in Liouville space. Details of this method can be found in the Appendix A.
Figure 1: Diffusion constant $D$ versus perturbation strength $\Delta^{\prime}$ for the model $H_{\text{NNN}}$ in Eq. (2) with anisotropy $\Delta=1.5$, as calculated numerically by dynamical quantum typicality for different system sizes NN and by the recursion method in the thermodynamic limit [48]. (a) $\langle S^{z}\rangle=0$, $tJ_{\text{max}}=10$. (b) $\langle S^{z}\rangle=0$, $tJ_{\text{max}}=40$. (c) $S^{z}=0$, $tJ_{\text{max}}=10$. (d) $S^{z}=0$, $tJ_{\text{max}}=40$. In (d), existing results from the microcanonical Lanczos method are also depicted [42].
@@
with non-negative rates $\alpha_{j}$, Lindblad operators $L_{j}$, and the anticommutator $\{\bullet,\bullet\}$. The Lindblad equation is the most general form of a quantum master equation, which is local in time and maps a density matrix to a density matrix again [49]. In the specific context of nonequilibrium transport, a common choice for the Lindblad operators is given by [6]
@@
Figure 3: (a) Diffusion constant $D$ versus perturbation strength $\Delta^{\prime}$ for the model $H_{\text{NNN}}$ in Eq. (2) with anisotropy $\Delta=1.5$, as calculated numerically in the open system for different system sizes NN. A comparison to the closed-system results from the recursion method is also indicated. (b) Extrapolation to the thermodynamic limit based on the linear scaling in (c).
A practical advantage of the Lindblad equation is that it can be solved by sophisticated numerical techniques, like stochastic unraveling [50, 15] or simulations based on matrix product states (MPS) [51, 52, 53]. In particular, MPS simulations are feasible for quite large system sizes. Yet, they are restricted to open boundary conditions and to large values of the system-bath coupling $\gamma$, due to the required reduction of entanglement growth. Thus, we use as an alternative strategy a recently proposed approach [54, 55, 56]. On the one hand, this approach is suitable for periodic boundary conditions and small system-bath coupling $\gamma$ [57]. On the other hand, it allows us to predict the open-system dynamics just on the knowledge of spatio-temporal correlation functions in the closed systems, which in turn enables us to treat open systems outside the range of exact diagonalization and stochastic unraveling. Details on this approach can be found in the Appendix B.
III Results
We now move on to our results, starting with the closed-system scenario and the model in Eq. (2). We first focus on results from the recursion method, which are, as discussed in the Appendix A, well converged w.r.t. to the number of Lanczos coefficients and depicted in Fig. 1 (a). As clearly visible, the diffusion constant $D$ has a continuous dependence on the perturbation strength $\Delta^{\prime}$, with no obvious signature of a discontinuity in the weak-perturbation limit $\Delta^{\prime}\to 0$. Furthermore, for the values of $\Delta^{\prime}$ considered, $D(\Delta^{\prime})$ has a minor dependence on $\Delta^{\prime}$ and is an almost flat curve.
In addition, we compare in Fig. 1 (a) the results from the recursion method to finite-size results for the quantity in Eq. (6) from dynamical quantum typicality, for which raw data can be found in the Appendix A. We do so for the grandcanonical ensemble $\langle S^{z}\rangle=0$ and at a time $tJ=10$. Apparently, the finite-size data is in very good agreement with the recursion method and shows no significant dependence on system size. When redoing the comparison for a longer time $tJ=40$ in Fig. 1 (b), the agreement becomes worse, due to finite-size effects at nonweak perturbation strengths $\Delta^{\prime}$. Still, when the system size is increased, the finite-size data gets closer and closer to the result from the recursion method, and they might eventually coincide in the thermodynamic limit.
@@ III Results
IV Simple model
To understand this observation, let us for a moment come back to the closed system and use a simple model for the quantity $D(t)$ in Eq. (6), as shown in Fig. 4. After an initial linear increase, $D(t)$ has a plateau but, after a finite-size time $\propto N$, increases linearly with a finite-size slope $\propto 1/N$. This overall behavior describes well the unperturbed case $\Delta^{\prime}=0$ with periodic boundary conditions [58]. It is worth mentioning that the linear slope is the finite-size Drude weight [6]. Using this model for $D(t)$, we can calculate the spreading of an excitation,
@@ IV Simple model
Assuming that $t^{*}$ is the relevant time scale for the open system, we plug in $t^{*}$ in the original model for $D(t)$ and use $D(t^{*})$ as an prediction for the diffusion constant in the open system. In Fig. 4 (b), we depict the finite-size scaling of the so obtained $D(t^{*})$, which is, despite the simplicity of arguments, very similar to the one for the open-system diffusion constant in Fig. 3 (c). Importantly, however, $D(t^{*})$ departs from a $1/N$-scaling for large $N$ and eventually diverges in the limit $N\to\infty$. The departure results from the fact that the steady-state time $t^{*}$ increases faster with $N$ than the time at which $D(t)$ is converged w.r.t. system size, see Appendix C. This finding opens the possibility of a diverging diffusion constant in the unperturbed open system with periodic boundary conditions. Additional results for open boundary conditions can be found in the Appendix B.
V Conclusion
We have analyzed the scaling of diffusion constants under integrability-breaking perturbations of the spin-$1/2$ XXZ chain in the easy-axis regime. To this end, we have used a variety of methods and covered both, closed and open systems, focusing on periodic boundary conditions. In the closed system, we have found evidence for a continuous change of diffusion constants over the full range of perturbation strengths. In the open system, we have found diffusion constants in quantitative agreement with the ones in the closed system in a range of nonweak perturbations. In the weak-perturbation limit, however, we have found disagreement and also pointed out the possibility of a diverging diffusion constant in the open system.
Acknowledgments
We thank Jochen Gemmer, Jacek Herbrych, Zala Lenarčič, Marcin Mierzejewski, and Peter Prelovšek for fruitful discussions.
@@ Acknowledgments
Appendix A Calculation of equilibrium transport
Figure 5: Diffusion constant $D_{R}$ versus $R$, as obtained from the recursion method for the model (a) $H_{\text{NNN}}$ in Eq. (2) and (b) $H_{\text{B}}$ in Eq. (3).
A.1 Recursion method
@@
${F}(s)=\int_{0}^{\infty}\text{d}t\,e^{ts}\,\tilde{C}(t)$ (13)
@@
$\tilde{C}(t)=\langle\tilde{j}(t)\tilde{j}\rangle\,,\quad\tilde{j}=\frac{j}{\sqrt{\langle j|j\rangle}}$ (14)
of the normalized current. In particular, ${\cal F}(0)$ determines the diffusion constant $D=\langle j^{2}\rangle F(0)/\chi$. According to the Mori formalism, Eq. (13) can be written in the form
${F}(s)=\frac{1}{s+\frac{b_{1}^{2}}{s+\frac{b_{2}^{2}}{s+\frac{b_{3}^{2}}{\cdots}}}}\,$, (15)
@@
$D=\frac{\langle j^{2}\rangle}{\chi b_{1}}\prod_{n=1}^{\infty}\left(\frac{b_{n}}{b_{n+1}}\right)^{(-1)^{n}}\,$, (16)
@@
It is now convenient to switch to the Hilbert space of operators and denote its elements $O$ by states $|O)$. This space is equipped with an inner product,
$(O_{m}|O_{n})=\text{tr}[O^{\dagger}_{m}O_{n}]\,$, (17)
which defines a norm via
$O||=\frac{\sqrt{(O|O)}}{\text{tr}[1]}\,.$ (18)
@@
Figure 6: Decay of the current autocorrelation function $\langle j(t)j\rangle$ in the unperturbed model in Eq. (1) with anisotropy $\Delta=1.5$ for (a) the grandcanonical ensemble $\langle S^{z}\rangle=0$ and (c) the canonical ensemble $S^{z}=0$, which is numerically calculated for different system sizes NN by the use of dynamical quantum typicality. (b) and (d) Corresponding time dependence of the diffusion coefficients $D(t)$.
Figure 7: The same data as in Fig. 6 but now for the model in Eq. (2) with $\Delta^{\prime}=0.4$.
@@
$b_{n}=||O^{\prime}_{n}||\,,\quad|O_{n})=\frac{|O^{\prime}_{n})}{b_{n}}\,.$ (22)
@@
$b_{n}=\alpha n+\beta\,,\quad n\geq R\,$, (23)
@@
$F_{0}\approx D_{R}=\begin{cases}\frac{1}{\tilde{p}_{R}b_{R}}\prod_{m=1}^{\frac{R}{2}}\frac{b_{2m}^{2}}{b_{2m-1}^{2}}&\text{, even }R\\ \frac{\tilde{p}_{R}}{b_{R}}\prod_{m=1}^{\frac{R-1}{2}}\frac{b_{2m}^{2}}{b_{2m-1}^{2}}&\text{, odd }R\end{cases}\,$, (24)
@@
$p_{R}=\frac{\Gamma(\frac{R}{2}+\frac{\beta}{2\alpha})\Gamma(\frac{R}{2}+\frac{\beta}{2\alpha}+1)}{\Gamma^{2}(\frac{R}{2}+\frac{\beta}{2\alpha}+\frac{1}{2})}\,.$ (25)
@@
$\alpha_{R}=b_{R}-b_{R-1}\,,\quad\beta_{R}=Rb_{R-1}-(R-1)b_{R}\,.$ (26)
In Fig. 5, we summarize the so obtained $D_{R}$. As mentioned in the main text, $D_{R}$ is well converged w.r.t. $R$, for the two models and the perturbation strengths considered.
A.2 Dynamical quantum typicality
@@
$\langle j(t)j\rangle=\frac{\langle\psi(t)|j|\varphi(t)\rangle}{\langle\psi(0)|\psi(0)\rangle}+{\cal O}\left(\frac{1}{\sqrt{d}}\right)\,$, (27)
@@
$\psi(t)=e^{-iHt}\,|\phi\rangle\,,\quad\varphi(t)=e^{-iHt}j\,|\phi\rangle\,$, (28)
and $|\phi\rangle$ is a Haar-random pure state. Importantly, the time argument in the approximation in Eq. (27) is now a property of the two pure states and not of the current operator anymore. These pure states can be propagated in time using fourth-order Runge-Kutta or Chebyshev polynomials, which then gives access to comparatively large system sizes outside the range of standard exact diagonalization. An approximation similar to the one in Eq. (27) can be also obtained for other operators, which we employ for the spatial-temporal correlation functions discussed later in the context of open systems.
In Figs. 6 - 8, we depict raw data for the current autocorrelation function $\langle j(t)j\rangle$ and its time integral $D(t)$ for three examples: The integrable model in Eq. (1) and the nonintegrable models in Eqs. (2) [with $\Delta^{\prime}=0.4$] and (3) [with $B/J=0.3$]. We do so for the two ensembles, grandcanonical $\langle S^{z}\rangle=0$ and canonical $S^{z}=0$. This and similar raw data is the basis of our analysis of diffusion constants in the main text. It is worth mentioning that the finite-size effects are unrelated to the approximation error in Eq. (27), which is negligibly small in all cases considered.
In Figs. 9 and 10, we additionally depict for the two nonintegrable cases the current autocorrelation function in frequency space, as obtained from a Fourier transform up to a maximum time $T_{\text{max}}$. As can be seen, the specific choice of $T_{\text{max}}$ controls the frequency resolution, and a reasonable choice in a finite system is not obvious. As an additional consistency check of our approach, we compare to existing data from the microcanonical Lanzos method [42] and find a convincing agreement, as visible in Figs. 9 (c) and 10 (d). MCLM data at zero frequency is also shown in the main text.
Figure 8: The same data as in Fig. 6 but now for the model in Eq. (3) with $B/J=0.3$.
Appendix B Calculation of nonequilibrium transport
B.1 Time evolution of densities
Figure 9: Diffusion coefficient $D$ versus frequency $\omega$ for the model in Eq. (2) with $\Delta^{\prime}=0.4$, as obtained numerically for the grandcanonical ensemble $\langle S^{z}\rangle=0$ using (a) fixed system size $N=24$ and different times $t_{\text{max}}$ and (b) fixed $t_{\text{max}}$ and different NN. (c) and (d) show (a) and (b) for the canonical ensemble $S^{z}=0$. In (c), the MCLM result from Ref. [42] is also depicted.
Figure 10: The same data as in Fig. 9 but now for the model in Eq. (3) with $B/J=0.3$.